# Corresponding Authors

### v1

@@ 3.3 Intermediate Motion Modeling with Diffusion
With a concrete synthesis network prepared, we move on to focus on modeling the intermediate motions for VFI. We use our fine-tuned flow estimator $\mathcal{F}$ as a teacher to train our motion diffusion model $\mathcal{M}$, which generates bi-directional optical flows from given two frames $I_{0}$ and $I_{1}$. Denoting concatenated flows $z_{0}=\{F_{\tau\rightarrow 0},F_{\tau\rightarrow 1}\}$, we train a diffusion model $\mathcal{M}$ by minimizing the following objective:
@@
where $z_{t}$ represents noisy flows diffused by Eq. 1. We concatenate $I_{0}$ and $I_{1}$ to $z_{t}$ and keep the teacher $\mathcal{F}$ frozen during training of $\mathcal{M}$. While $\epsilon$-prediction [18] and $L_{2}$ norm are popular choices for training diffusion-based image generative models, we found $x_{0}$-prediction and $L_{1}$ norm to be beneficial for modeling flows. While image diffusion models utilize a U-Net architecture that employs input and output of the same resolution for noisy images that operates fully on the entire resolution, we introduce a new architecture for $\mathcal{M}$ aimed at learning optical flows and enhancing efficiency, which will be described in the following paragraph.
Figure 3: Visualization of our diffusion model architecture. The inputs of two frames and two noisy flow maps, are downsampled and projected respectively, to an 8×8$\times$8 × smaller resolution. It then goes through a 3-level U-Net, which outputs a pair of coarse flow maps and their corresponding weight masks for upsampling. The convex upsampling layer takes coarse flow maps and the weight masks as the input, and upsamples the coarse flow to the original resolution. Note that the convex upsampling layer accordingly do not have any learnable parameters.
3.3.1 Architecture
In our novel diffusion model architecture designed for motion modeling, we begin by excluding attention layers, as we have found doing so saves memory without deterioration in performance. Observing that optical flow maps—our primary target—are sparse representations encoding only low-frequency information, we opt to avoid the unnecessary complexity of full-resolution flow estimation. Consequently, we predict flows at 1/8 the original resolution, a method mirroring the coarse-to-fine strategies [69, 66, 65, 21], thus sidestepping the need for full-resolution flow estimation. We realize this by introducing input downsampling and convex upsampling, making our architecture computationally efficient and well-suited to meet our resolution-specific needs. We elaborate them in the following paragraphs.
@@ 3.3.1 Architecture
Given an input $\{z_{t},~{}I_{0},~{}I_{1}\}$ of 10 channels, we downsample it to $1/8$ size. Rather than using a single layer to directly apply on the 10 channel input, we separately apply different layers on the frames and the noisy flows:
@@
Once we obtain the downsampled features $I_{0}^{\prime},I_{1}^{\prime},z_{t}^{\prime}$, we concatenate and project them to features:
@@
where the projection layer $\mathcal{D}_{p}$ is implemented by a single $1\times 1$ convolutional layer. By sharing the parameters applied to $I_{0}$ and $I_{1}$, we could save the number of parameters required for the downsampling process. In addition, it makes it invariant to the order of taking the two frames.
@@
Starting with a coarse flow estimation of size $H/8\times W/8$ from an initial $H\times W$ input, we upsample the predicted flow to the original resolution using a weighted combination of a $3\times 3$ grid from the coarse flow. Utilizing skip connections from the U-Net encoder and a residual block, we generate a mask with $8\times 8\times 9$ channels. Applying softmax to calculate the weights for the 9 neighboring pixels, we then perform a weighted summation to produce the refined flow map. This method aligns with the $\textbf{x}_{0}$-prediction, promoting local correlations and differing from $\epsilon$-predictions which necessitate locally independent estimations.
By operating at a reduced resolution, specifically at 64×64$\times$64 × smaller space, we achieve significant computation savings. This approach of using diffusion models at a compressed resolution resembles the strategy of latent diffusion models [54], which are noted for their efficiency. Note that this convex upsampling layer does not involve any learnable parameters.
@@ 4.1 Experiment Settings
4.1.1 Implementation Details
We train our model on the Vimeo90k dataset [71], comprising 51,312 triplets designed for VFI, using random $256\times 256$ crops with augmentations like 90 rotation, flipping, and frame order reversing. Since our motion diffusion model is trained on a well-curated data ranging within 256 resolution, it could suffer from a performance drop when it comes to high resolution videos of large motions which goes beyond the distribution of the training data. To handle these cases, we generate flows at the training resolution by resizing the inputs, followed by post-processing of bicubic upsampling at inference time. We recommend the reader to refer to the Appendix for further details.
4.1.2 Evaluation Protocol
We evaluate on well-known VFI benchmarks: SNU-FILM [6], Middlebury (others-set) [1], and Xiph [41, 45], chosen for their broad motion diversity and magnitudes. SNU-FILM is divided into 4 subsets (easy, medium, hard, and extreme) with 310 triplets each, totaling 1240 triplets. The Middlebury benchmark, particularly the others-set, includes 12 frame triplets with a resolution of approximately $640\times 480$. The Xiph benchmark features two subsets, Xiph-2K and Xiph-4K, with 392 frame triplets in 4K resolution. The 2K subset contains 4K triplets downsampled to 2K, while the 4K subset consists of center-cropped 2K patches from the 4K triplets, preserving the original motion magnitudes.
Following practices in generative models-based restoration [8, 33, 46], we use perceptual metrics, LPIPS [73] and DISTS [10] for evaluation. While PSNR and SSIM are popular metrics in evaluating reconstruction quality, they have been known to differ from human perception in some aspects, sensitive to imperceptible differences in pixels and preferring blurry samples [73]. Since our method is based on generative models, we prioritize LPIPS and DISTS, which highly correlates with human perception. The full results including the fidelity metrics such as PSNR and SSIM can be found in the Appendix.
Method SNU-FILM-easy SNU-FILM-medium SNU-FILM-hard SNU-FILM-extreme
@@ 4.1.2 Evaluation Protocol
XVFI [61] 0.0251 0.0219 0.0389 0.0304 0.0668 0.0431 0.1189 0.0638
XVFIv [61] 0.0175 0.0181 0.0322 0.0276 0.0629 0.0414 0.1257 0.0673
RIFE [22] 0.0181 0.0195 0.0317 0.0289 0.0657 0.0443 0.1390 0.0764
IFRNet [28] 0.0201 0.0206 0.0320 0.0281 0.0573 0.0398 0.1176 0.0653
FILM-$\mathcal{L}_{1}$ [52] 0.0184 0.0217 0.0315 0.0316 0.0568 0.0441 0.1060 0.0632
ABME [49] 0.0222 0.0229 0.0372 0.0344 0.0658 0.0496 0.1258 0.0747
IFRNet-Large [28] 0.0203 0.0211 0.0321 0.0288 0.0562 0.0403 0.1131 0.0638
UPRNet [24] 0.0179 0.0200 0.0322 0.0311 0.0604 0.0452 0.1115 0.0653
AMT-G [32] 0.0325 0.0312 0.0447 0.0395 0.0680 0.0506 0.1128 0.0686
EMA-VFI [72] 0.0186 0.0204 0.0325 0.0318 0.0579 0.0457 0.1099 0.0671
UPRNet-LARGE [24] 0.0182 0.0203 0.0334 0.0327 0.0612 0.0475 0.1109 0.0672
CAIN [6] 0.0197 0.0229 0.0375 0.0347 0.0885 0.0606 0.1790 0.1042
FILM-$\mathcal{L}_{vgg}$ [52] 0.0123 0.0128 0.0219 0.0183 0.0443 0.0282 0.0917 0.0471
FILM-$\mathcal{L}_{s}$ [52] 0.0120 0.0124 0.0213 0.0177 0.0429 0.0268 0.0889 0.0448
LDMVFI [8] 0.0145 0.0130 0.0284 0.0219 0.0602 0.0379 0.1226 0.0651
@@
[…]
Method Middlebury Xiph-2K Xiph-4K
LPIPS DISTS LPIPS DISTS LPIPS DISTS
XVFI [61] 0.0520 0.0557 0.0710 0.0360 0.1222 0.0610
XVFIv [61] 0.0169 0.0244 0.0844 0.0418 0.1835 0.0779
RIFE [22] 0.0162 0.0228 0.0918 0.0481 0.2072 0.0915
FILM-$\mathcal{L}_{1}$ [52] 0.0173 0.0246 0.0906 0.0510 0.1841 0.0884
ABME [49] 0.0290 0.0325 0.1071 0.0581 0.2361 0.1108
IFRNet-Large [28] 0.0285 0.0366 0.0681 0.0372 0.1364 0.0665
AMT-G [32] 0.0486 0.0533 0.1061 0.0563 0.2054 0.1005
EMA-VFI [72] 0.0151 0.0218 0.1024 0.0550 0.2258 0.1049
UPRNet-LARGE [24] 0.0150 0.0209 0.1010 0.0553 0.2150 0.1017
CAIN [6] 0.0254 0.0383 0.1025 0.0533 0.2229 0.0980
FILM-${\mathcal{L}_{vgg}}^{{\dagger}}$ [52] 0.0096 0.0148 0.0355 0.0238 0.0754 0.0406
FILM-$\mathcal{L}_{s}$ [52] 0.0093 0.0140 0.0330 0.0237 0.0703 0.0385
LDMVFI [8] 0.0195 0.0261 0.0420 0.0163 0.0859 0.0359
MoMo (Ours) 0.0094 0.0126 0.0300 0.0119 0.0631 0.0274
[…]
@@ 4.2 Comparison to State-of-the-arts
To validate the effectiveness of our approach, we compare MoMo with state-of-the-art VFI methods. The methods for comparison include: CAIN [6], FILM-$\mathcal{L}_{vgg}$, FILM-$\mathcal{L}_{s}$ and LDMVFI [8], which employ with perceptual-oriented loss in the training process. Since models trained for improved perceptual quality are publicly available by a very limited number, we also include methods trained with the traditional pixel-wise reconstruction loss: XVFI [61], RIFE [22], IFRNet [28], FILM-$\mathcal{L}_{1}$ [52], ABME [49], UPRNet [24], AMT [32], EMA-VFI [72].
4.2.1 Quantitative Results
Tables 1 and 2 present our quantitative results across three benchmark datasets. We order the baseline methods which rely on the traditional pixel-wise reconstruction loss by their PSNR scores on SNU-FILM’s ‘hard’ subset, with lower and higher PSNR scores at the top and bottom, respectively. This order suggests PSNR may not always reflect human perception, with higher PSNR models exhibiting lower perceptual quality at times.
MoMo achieves state-of-the-art on all SNU-FILM subsets, leading in both LPIPS and DISTS metrics. On Middlebury, it outperforms baselines in DISTS and closely trails FILM-$\mathcal{L}_{s}$ in LPIPS. MoMo also excels on both Xiph subsets, 2K and 4K, in both metrics. MoMo surpasses methods like XVFI, FILM, and UPRNet, tailored for high-resolution flow estimation. This highlights the effectiveness of our approach in generating well-structured optical flows for the intermediate frame through proficient intermediate motion modeling.
@@ 4.2.1 Quantitative Results
Figure 4: Visualized comparison of estimated intermediate flows against state-of-the-art methods. Although XVFI [61] and FILM [52] is designed for estimating of elaborate flows on high resolution images, it sometimes fails to predict well-structured flow fields, leading to unpleasant synthesis results. On the other hand, our flow estimations consistently show a well-structured flow fields which leads to promising synthesis of frames.
4.2.2 Qualitative Results
The qualitative results of MoMo with comparison to the state-of-the-art algorithms can be found at Fig. 1 and 5. In Fig. 1, it can be found that MoMo reconstructs both wings with rich details, whereas methods of the top row, which greatly relies on the $L_{1}$ pixel-wise loss in training, namely ABME [49], IFRNet [28] and AMT [32] all show blurry results. Moreover, our result also outperforms state-of-the-art models designed particularly for perceptual quality, LDMVFI [8] and FILM-$\mathcal{L}_{s}$ [52], with well-structurized synthesis of both wings. Fig. 5 present additional results obtained from the ‘extreme’ subset of SNU-FILM and Xiph-4K set. Here, we compare with LDMVFI [8] and FILM-$\mathcal{L}_{s}$ [52], methods designed specifically for visual quality, along with FILM-$\mathcal{L}_{1}$ [52]. We use FILM-$\mathcal{L}_{1}$ [52] as a representative among the $L_{1}$ reconstruction-based method, as it performed the most robust and competently in terms of perceptual metrics across multiple benchmarks as reported in Tab. 1 and 2. MoMo consistently shows a superior visual quality, with less artifacts and well-structured objects.
@@ 4.3 Ablation Studies
To verify the effects of our design choices, the ablation studies in the following sections are conducted on the ‘hard’ subset of SNU-FILM dataset, unless mentioned otherwise. We start by studying the effects of the teacher optical flow model, used for training the motion diffusion model. We then experiment on the number of denoising steps used at inference time. Last but not least, we study on the design choices in diffusion architecture, namely the use of $\textbf{x}_{0}$-prediction and our U-Net with convex upsampling.
Teacher LPIPS DISTS
@@
Fine-tuned (proposed) 0.0419 0.0252
Table 3: Ablation on the teacher flow model. We use the RAFT [66] model fixed to the pre-trained state, fine-tuned after training the synthesis model, and trained with the synthesis model in an end-to-end manner. Our experiments show that fine-tuning after training the synthesis model is the most effective.
@@
Table 4: Experiment on the number of denoising steps for inference (on SNU-FILM-extreme). Our experiments show that about 8 steps is enough, and use of more steps exceeding this does not lead to a notable improvement considering the runtime tradeoff.
4.3.1 Optical Flow Teacher
We conduct ablation study on the teacher optical flow model. We choose RAFT [66] as the teacher model, which is the state-of-the-art architecture for optical flow estimation. We prepare the teacher model of three different weights. 1) we use the default off-the-shelf weights provided by the Torchvision library [37],trained specifically for optical flow estimation on multiple datasets [13, 38, 3, 40, 27]. 2) we fine-tune the pre-trained RAFT with our frozen synthesis model as described in Sec. 3.2. 3) we jointly train RAFT and our synthesis network in end-to-end manner, with RAFT initialized with the pre-trained weights. We use these three RAFT models of different weights as the teacher for the experiment.
The ablation study summarized in Table 4 shows that fine-tuning the flow model $\mathcal{F}$ after training the synthesis model $\mathcal{G}$ is the most effective, while concurrent training is the least effective. Fine-tuning of $\mathcal{F}$ enhances flow estimation and suitability for synthesis tasks [71]. However, end-to-end training can cause the synthesis model to depend too heavily on estimated flows, risking inaccuracies from the motion diffusion model. Our results highlight that sequential training of the synthesis model and the flow estimator ensures optimal performance.
4.3.2 Number of Steps
@@
Table 4 shows the effect of number of denoising steps in motion generation, experimented on the ‘extreme’ subset of SNU-FILM. We find that around 8 steps is optimal, with more steps not markedly improving performance. This is likely due to the simpler nature of flow representations compared to RGB pixels. In contrast to image diffusion models [54] and LDMVFI [8] which requires over 50 steps, our method delivers satisfactory outcomes with far fewer steps, cutting down on both runtime and computational expenses.
4.3.3 Diffusion Architecture
Data Prediction Architecture LPIPS DISTS FLOPs (T) Params. (M) Runtime (ms)
Latent $\epsilon$ Standard U-Net [8] 0.0601 0.0379 3.25 439.0 10283.51
Flow $\epsilon$ Standard U-Net 0.4090 0.2621 8.08 69.1 603.64
Flow $\textbf{x}_{0}$ Standard U-Net 0.0460 0.0295 8.08 69.1 603.64
Flow $\textbf{x}_{0}$ (weighted) Convex-Up U-Net (Ours) 0.0463 0.0298 1.12 71.6 145.49
Flow $\textbf{x}_{0}$ Convex-Up U-Net (Ours) 0.0419 0.0252 1.12 71.6 145.49
Table 5: We conduct an ablation study on our diffusion model, emphasizing its minimal computational needs and fastest runtime. The results showcase the entire framework’s performance, covering both diffusion and synthesis stages. For reference, the first row includes LDMVFI’s approach, which relies on a larger diffusion model and a vector-quantized [14] decoder, leading to a notably slower runtime than ours.
In our ablation study, detailed in Table 5, we assess our motion diffusion model using the standard timestep-conditioned U-Net architecture (UNet2DModel) from the diffusers library [50], alongside $\epsilon$- and $\textbf{x}_{0}$-prediction types. Contrary to the common preference for $\epsilon$-prediction in diffusion models, our motion diffusion model favors $\textbf{x}_{0}$-prediction.
We also explore our coarse-to-fine estimation using convex upsampling [66]. This approach reduces computational costs by operating at a lower resolution and improves performance. Given that our architecture predicts values with a strong correlation between neighboring pixels, $\epsilon$-prediction, which samples noise independently, proves less suitable. We experiment with a SNR-weighted $\textbf{x}_{0}$-prediction [59], to make it equivalent to $\epsilon$-prediction loss. Nonetheless, $\textbf{x}_{0}$-prediction consistently outperforms, validating our architectural decisions.
Despite having a similar number of parameters as the standard U-Net, our Convex Upsampling U-Net significantly reduces floating point operations (FLOPs) by about 7.2×7.2$\times$7.2 ×. Runtime tests on a NVIDIA 32GB V100 GPU for $256\times 448$ resolution frames—averaged over 100 iterations—reveal that our Convex-Up U-Net processes frames in approximately 145.49 ms each, achieving a 4.15×4.15$\times$4.15 × speedup over the standard U-Net and an 70×70$\times$70 × faster inference speed than the LDMVFI [8] baseline. This efficiency is attributed to our model’s efficient architecture and notably fewer denoising steps.
5 Limitation and Future Work
As mentioned in Sec. 4.1.1, our framework generates flows at a resized resolution, followed by bicubic upsampling during inference. Consequently, the generated flows could lack in details at motion boundaries, although it still performs superior to baseline methods. By adopting an recurrent architecture for the motion diffusion model, as in recent efforts in large motion estimation [61, 52, 24], we believe our approach could be further reinforced, and leave this for future work.

### v2

@@ 3.3 Intermediate Motion Modeling with Diffusion
With a synthesis network fixed, we focus on modeling the intermediate motions for VFI. We use our fine-tuned flow model $\mathcal{F}$ as the teacher to train our motion diffusion model $\mathcal{M}$, which generates bi-directional optical flows from the pair of input frames $I_{0}$ and $I_{1}$. Denoting concatenated flows $z_{0}=\{F_{\tau\rightarrow 0},F_{\tau\rightarrow 1}\}$, we train a diffusion model $\mathcal{M}$ by minimizing the following objective:
@@
where $z_{t}$ represents noisy flows diffused by Eq. 1. We concatenate $I_{0}$ and $I_{1}$ to $z_{t}$ and keep the teacher $\mathcal{F}$ frozen during training of $\mathcal{M}$. While $\epsilon$-prediction (Ho, Jain, and Abbeel 2020) and $L_{2}$ norm are popular choices for training diffusion-based image generative models, we found $\textbf{x}_{0}$-prediction and $L_{1}$ norm to be beneficial for modeling flows. While image diffusion models utilize a U-Net architecture that employs input and output of the same resolution for noisy images that operates fully on the entire resolution, we introduce a new architecture for $\mathcal{M}$ aimed at learning optical flows and enhancing efficiency, which will be described in the following paragraph.
Architecture
An overview of our proposed motion diffusion model architecture is provided at Fig. 3. In our novel diffusion model architecture designed for motion modeling, we begin by excluding attention layers, as we have found doing so saves memory without deterioration in performance. Observing that optical flow maps—our primary target—are sparse representations encoding low-frequency information, we opt to avoid the unnecessary complexity of full-resolution estimation. Consequently, we initially predict flows at $1/8$ of the input resolution and upsample them by 8×8$\times$8 ×, a method mirroring the coarse-to-fine strategies (Teed and Deng 2020; Xu et al. 2022; Huang et al. 2022a), thus sidestepping the need for full-resolution flow estimation. We realize this by introducing input downsampling and convex upsampling (Teed and Deng 2020) into U-Net, making our architecture computationally efficient and well-suited to meet our resolution-specific needs. We elaborate them in the following paragraphs.
@@
Given an input $\{z_{t},I_{0},I_{1}\}$ of 10 channels, we downsample and encode it to $1/8$ resolution. Rather than using a single layer to directly apply on the 10 channel input, we separately apply layers $\mathcal{D}_{I}$ and $\mathcal{D}_{z}$ on the frames and the noisy flows, respectively:
@@
By sharing the parameters applied to $I_{0}$ and $I_{1}$, we could save the number of parameters required for the downsampling process, and make it invariant to the order of the two input frames. Once we obtain the downsampled features $I_{0}^{\prime},I_{1}^{\prime},z_{t}^{\prime}$, we concatenate and project them to features:
@@
where the projection layer $\mathcal{D}_{p}$ is implemented by a single $1\times 1$ convolutional layer.
The projected features $p_{t}$ is then given to a 3-level diffusion U-Net, which produces two outputs: a coarse estimation of flow maps and their corresponding upsampling weight masks. The two outputs are combined in the convex upsampling layer to obtain the final full-scale flow maps.
@@
Given the coarse flow maps and their corresponding upsampling weight masks, both of size $H/8\times W/8$, we attempt to upsample the coarse flows to the original $H\times W$ resolution using a weighted combination of $3\times 3$ grid of each coarse flow neighbors, by integrating the convex upsampling layer (Teed and Deng 2020) to our architecture. Using the predicted upsampling weight masks of $8\times 8\times 9$ channels, we apply softmax on the weights of 9 neighboring pixels, and perform weighted summation with coarse flows to obtain the final upsampled flow map. An illustrated description is provided in the Appendix.
This method aligns with $\textbf{x}_{0}$-prediction, promoting local correlations and differing from $\epsilon$-predictions which necessitate locally independent estimations. By operating at a reduced resolution, specifically at 64×64$\times$64 × smaller space, we achieve significant computation savings. Note that this upsampling layer does not involve any learnable parameters.
@@ 4.1 Experiment Settings
Implementation Details
We train our model on the Vimeo90k dataset (Xue et al. 2019), using random $256\times 256$ crops with augmentations like 90 rotation, flipping, and frame order reversing. We recommend the reader to refer to the Appendix for further details.
Evaluation Protocol
We evaluate on well-known VFI benchmarks: Vimeo90k (Xue et al. 2019), SNU-FILM (Choi et al. 2020), Middlebury (others-set) (Baker et al. 2011), and Xiph (Montgomery and Lars 1994; Niklaus and Liu 2020), chosen for their broad motion diversity and magnitudes. Following practices in generative models-based restoration (Danier, Zhang, and Bull 2024; Liang, Zeng, and Zhang 2022), we focus on perceptual similarity metrics LPIPS (Zhang et al. 2018) and DISTS (Ding et al. 2020), which highly correlates with human perception, for evaluation. While PSNR and SSIM are popular metrics, they have been known to differ from human perception in some aspects, sensitive to imperceptible differences in pixels and preferring blurry samples (Zhang et al. 2018). The full results can be found in the Appendix.
Method Perception- oriented loss FILM-easy FILM-medium FILM-hard FILM-extreme
@@
ABME (Park, Lee, and Kim 2021) 0.0222 0.0229 0.0372 0.0344 0.0658 0.0496 0.1258 0.0747
XVFIv (Sim, Oh, and Kim 2021) 0.0175 0.0181 0.0322 0.0276 0.0629 0.0414 0.1257 0.0673
IFRNet-Large (Kong et al. 2022) 0.0203 0.0211 0.0321 0.0288 0.0562 0.0403 0.1131 0.0638
RIFE (Huang et al. 2022b) 0.0181 0.0195 0.0317 0.0289 0.0657 0.0443 0.1390 0.0764
FILM-$\mathcal{L}_{1}$ (Reda et al. 2022) 0.0184 0.0217 0.0315 0.0316 0.0568 0.0441 0.1060 0.0632
AMT-G (Li et al. 2023) 0.0325 0.0312 0.0447 0.0395 0.0680 0.0506 0.1128 0.0686
EMA-VFI (Zhang et al. 2023) 0.0186 0.0204 0.0325 0.0318 0.0579 0.0457 0.1099 0.0671
UPRNet-LARGE (Jin et al. 2023) 0.0182 0.0203 0.0334 0.0327 0.0612 0.0475 0.1109 0.0672
CAIN (Choi et al. 2020) 0.0197 0.0229 0.0375 0.0347 0.0885 0.0606 0.1790 0.1042
FILM-$\mathcal{L}_{vgg}$ (Reda et al. 2022) 0.0123 0.0128 0.0219 0.0183 0.0443 0.0282 0.0917 0.0471
FILM-$\mathcal{L}_{s}$ (Reda et al. 2022) 0.0120 0.0124 0.0213 0.0177 0.0429 0.0268 0.0889 0.0448
LDMVFI (Danier, Zhang, and Bull 2024) 0.0145 0.0130 0.0284 0.0219 0.0602 0.0379 0.1226 0.0651
PerVFI (Wu et al. 2024) 0.0142 0.0124 0.0245 0.0181 0.0561 0.0635 0.0902 0.0448
@@
Table 1: Quantitative experiments on the SNU-FILM benchmark (Choi et al. 2020). The best results are in bold, and the second best is underlined, respectively. Our method outperforms existing methods on all four subsets.
Method Perception- oriented loss Middlebury Vimeo90k Xiph-2K Xiph-4K
LPIPS DISTS LPIPS DISTS LPIPS DISTS LPIPS DISTS
ABME (Park, Lee, and Kim 2021) 0.0290 0.0325 0.0213 0.0353 0.1071 0.0581 0.2361 0.1108
XVFIv (Sim, Oh, and Kim 2021) 0.0169 0.0244 0.0229 0.0354 0.0844 0.0418 0.1835 0.0779
IFRNet-Large (Kong et al. 2022) 0.0285 0.0366 0.0189 0.0325 0.0681 0.0372 0.1364 0.0665
RIFE (Huang et al. 2022b) 0.0162 0.0228 0.0223 0.0356 0.0918 0.0481 0.2072 0.0915
FILM-$\mathcal{L}_{1}$ (Reda et al. 2022) 0.0173 0.0246 0.0197 0.0343 0.0906 0.0510 0.1841 0.0884
AMT-G (Li et al. 2023) 0.0486 0.0533 0.0195 0.0351 0.1061 0.0563 0.2054 0.1005
EMA-VFI (Zhang et al. 2023) 0.0151 0.0218 0.0196 0.0343 0.1024 0.0550 0.2258 0.1049
UPRNet-LARGE (Jin et al. 2023) 0.0150 0.0209 0.0201 0.0342 0.1010 0.0553 0.2150 0.1017
CAIN (Choi et al. 2020) 0.0254 0.0383 0.0306 0.0483 0.1025 0.0533 0.2229 0.0980
FILM-${\mathcal{L}_{vgg}}$ (Reda et al. 2022) 0.0096 0.0148 0.0137 0.0229 0.0355 0.0238 0.0754 0.0406
FILM-$\mathcal{L}_{s}$ (Reda et al. 2022) 0.0093 0.0140 0.0131 0.0224 0.0330 0.0237 0.0703 0.0385
LDMVFI (Danier, Zhang, and Bull 2024) 0.0195 0.0261 0.0233 0.0327 0.0420 0.0163 0.0859 0.0359
PerVFI (Wu et al. 2024) 0.0142 0.0163 0.0179 0.0248 0.0381 0.0153 0.0858 0.0331
MoMo (Ours) 0.0094 0.0126 0.0136 0.0203 0.0300 0.0119 0.0631 0.0274
Table 2: Quantitative experiments on the three benchmarks, Middlebury (Baker et al. 2011), Vimeo90k (Xue et al. 2019) and Xiph-2K,4K (Montgomery and Lars 1994; Niklaus and Liu 2020). The best results are in bold, and the second best is underlined, respectively.
@@ 4.2 Comparison to State-of-the-arts
Baselines
We compare our method, MoMo, with state-of-the-art VFI methods which employ perception-oriented objectives in the training process: CAIN (Choi et al. 2020), FILM-$\mathcal{L}_{vgg}$, FILM-$\mathcal{L}_{s}$ (Reda et al. 2022), LDMVFI (Danier, Zhang, and Bull 2024) and PerVFI (Wu et al. 2024). Since there is a limited number of methods focused on perception-oriented objectives, we also include methods trained with the traditional pixel-wise reconstruction loss for comparison: XVFI (Sim, Oh, and Kim 2021), RIFE (Huang et al. 2022b), IFRNet (Kong et al. 2022), AMT (Li et al. 2023), EMA-VFI (Zhang et al. 2023), UPRNet (Jin et al. 2023).
Quantitative Results
Tables 1 and 2 present our quantitative results across four benchmark datasets. MoMo achieves state-of-the-art on all four subsets of SNU-FILM, leading in both LPIPS and DISTS metrics. On Middlebury and Vimeo90k, it outperforms baselines in DISTS and closely trails FILM-$\mathcal{L}_{s}$ in LPIPS. MoMo also excels on both Xiph subsets, 2K and 4K, in both metrics. This highlights the effectiveness of our approach in generating well-structured optical flows for the intermediate frame through proficient intermediate motion modeling.
@@
Figure 4: Visualized comparison of estimated intermediate flows against state-of-the-art methods. Our flow estimations show better-structured flow fields which leads to promising synthesis of frames.
@@
Qualitative Results
The qualitative results of MoMo with comparison to the state-of-the-art algorithms can be found at Fig. 1 and 5. In Fig. 1, MoMo reconstructs both wings with rich details, whereas methods of the top row, which greatly relies on the $L_{1}$ pixel-wise loss in training show blurry results. Moreover, our result also outperforms state-of-the-art models designed particularly for perceptual quality, LDMVFI and FILM-$\mathcal{L}_{s}$, with well-structurized synthesis of both wings. Fig. 5 present additional results obtained from the ‘extreme’ subset of SNU-FILM and Xiph-4K set. MoMo consistently shows a superior visual quality, with less artifacts and well-structured objects. The interpolated video samples are provided in the supplementary material.
@@ 4.3 Ablation Studies
We conduct ablation studies to verify the effects of our design choices. We use the ‘hard’ subset of SNU-FILM dataset, unless mentioned otherwise. We start by studying the effects of the teacher optical flow model, used for training the motion diffusion model. We then experiment on the number of denoising steps used at inference time. Lastly, we study on the design choices in diffusion architecture.
Data Prediction Architecture LPIPS DISTS TFLOPs Params. (M) R-time (ms)
Latent $\epsilon$ Standard U-Net (Danier, Zhang, and Bull 2024) 0.0601 0.0379 3.25 439.0 10283.51
Flow $\epsilon$ Standard U-Net 0.4090 0.2621 8.08 71.1 603.64
Flow $\textbf{x}_{0}$ Standard U-Net 0.0460 0.0295 8.08 71.1 603.64
Flow $\textbf{x}_{0}$ (weighted) Convex-Up U-Net (Ours) 0.0463 0.0298 1.12 73.6 145.49
Flow $\textbf{x}_{0}$ Convex-Up U-Net (Ours) 0.0425 0.0257 1.12 73.6 145.49
Flow $\textbf{x}_{0}$ Convex-Up U-Net (Ours) + Longer Training 0.0419 0.0252 1.12 73.6 145.49
Table 3: Ablation study on our motion diffusion model. Our design choice reaches the best performance with minimal computational needs and fastest runtime. The first row includes our baseline, LDMVFI (Danier, Zhang, and Bull 2024), for reference.
Optical Flow Teacher
We conduct a study on the teacher optical flow model. We choose RAFT (Teed and Deng 2020) as the teacher model, the state-of-the-art model for optical flow estimation. We test the teacher model of three different weights: 1) the default off-the-shelf weights provided by the Torchvision library (maintainers and contributors 2016). 2) Initialized with the pre-trained weights, weights trained jointly with our synthesis network in an end-to-end manner. 3) Optimized in an alternating manner with the synthesis network, initialized from pre-trained weights, as described in Sec. 3.2. We use these three different versions of RAFT as the teacher for the experiment.
The ablation study summarized in Table 4 shows that fine-tuning the flow model $\mathcal{F}$ after training the synthesis model $\mathcal{G}$ is the most effective. Fine-tuning of $\mathcal{F}$ enhances flow estimation and suitability for synthesis tasks (Xue et al. 2019). However, end-to-end training can cause the synthesis model to depend too heavily on estimated flows, risking inaccuracies from the motion diffusion model. Our results highlight that sequential training of the synthesis model and the flow estimator ensures optimal performance.
Diffusion Architecture
In our ablation study, detailed in Table 3, we assess our motion diffusion model using the standard timestep-conditioned U-Net architecture (UNet2DModel) from the diffusers library (von Platen et al. 2022), alongside $\epsilon$- and $\textbf{x}_{0}$-prediction types. Contrary to the common preference for $\epsilon$-prediction in diffusion models, our motion diffusion model favors $\textbf{x}_{0}$-prediction.
We also explore our coarse-to-fine estimation using convex upsampling. This approach reduces computational costs and improves performance. Given that our architecture predicts values with a strong correlation between neighboring pixels, $\epsilon$-prediction, which samples noise independently, proves less suitable. We experiment with a SNR-weighted $\textbf{x}_{0}$-prediction (Salimans and Ho 2022), to make it equivalent to $\epsilon$-prediction loss. Nonetheless, $\textbf{x}_{0}$-prediction consistently outperforms, validating our architectural decisions.
Despite having a similar number of parameters as the standard U-Net, our Convex Upsampling U-Net significantly reduces floating point operations (FLOPs) by about 7.2×7.2$\times$7.2 ×. Runtime tests on a NVIDIA 32GB V100 GPU for $256\times 448$ resolution frames—averaged over 100 iterations—reveal that our Convex-Up U-Net processes frames in approximately 145.49 ms each, achieving a 4.15×4.15$\times$4.15 × speedup over the standard U-Net and an 70×70$\times$70 × faster inference speed than the LDMVFI baseline. This efficiency is attributed to our model’s efficient architecture and notably fewer denoising steps.
Teacher Fine-tune $\mathcal{F}$ Train $\mathcal{S}$ LPIPS DISTS
@@
Alternating (Ours) 0.0419 0.0252
Table 4: Experiments on the teacher flow model. We use RAFT (Teed and Deng 2020) with three different weights. The results show that alternating optimization, fine-tuning the flow model with $\mathcal{S}$ fixed, to be the most effective.
@@
Table 5: Experiment on the number of denoising steps for inference (on SNU-FILM-extreme). Our experiments show that about 8 steps is enough, and use of more steps exceeding this does not lead to a notable improvement.
Effectiveness of Diffusion
@@
Table 5 shows the effect of number of denoising steps in motion generation, experimented on the ‘extreme’ subset of SNU-FILM. We observe consistent improvement with more number of steps up to 8, with more steps not markedly improving performance. In contrast to image diffusion models (Rombach et al. 2022) and LDMVFI (Danier, Zhang, and Bull 2024) which requires over 50 steps, our method delivers satisfactory outcomes with far fewer steps, cutting down on both runtime and computational expenses. This is likely due to the simpler nature of flow representations compared to RGB pixels. This experiment shows the effectiveness of using diffusion models in motion modeling, as use of multiple steps guarantees better motion predictions.